OpenStudy (anonymous):

Please help me to find the slope of the tangent line to the given curve at the given value of x and find the equation of each tangent line : Y=8-x^2 at x=1 I don't just need a answer, I need to know how to work it....Thanks in advance :)

OpenStudy (anonymous):

see differentiate wrt x on both sides

OpenStudy (anonymous):

both sides of what?

OpenStudy (anonymous):

both sides of ur equation y = 8 - x^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do it and show it to me

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

sorry computer froze...1 sec

OpenStudy (anonymous):

I thought that in order to find the slope of a tangent it would be to write f(x) = 8-x^2 and a=1

OpenStudy (anonymous):

wts a?

OpenStudy (anonymous):

Slope of a tangent should be ((1+h)^2+8) - ...............disregard.....Thank You!

OpenStudy (anonymous):

youre going totally off track now differentiate both sides dy/dx = -2x slope is dy/dx at x=1 hence slope is -2 get it?

OpenStudy (anonymous):

the equation of the tangent line should be \[\lim_{h \rightarrow 0} [f \left( x+h \right)-f \left( x \right)]\div h\]which will get you -2x and by placing x=1 you can find the slope at x=1

OpenStudy (anonymous):

do u know differentiation?

OpenStudy (anonymous):

I apologize but my computer is acting up tonight........

OpenStudy (anonymous):

@ kanade, that is the formula I was using, that is why I was so confused with @him1618, I am not that good in calculus/stats so if you tell me terms and stuff I get lost

OpenStudy (anonymous):

find a tangent gradient with diferentiate y=8-x^2 to y'=-2x fill x with the point x=1 and then you will get the gradient, use it to get equation with formula y-y1=m(x-x1), which m = gradient, (x1,y1) is a point which we want to know about equation